1 The Law of large numbers implies that the relative frequency of an event after many trials is increasingly predictable; this however cannot predict the outcome of any individual trial True False

2 In a batch of 25 products, there are 9 defective units. 6 of these have only minor defects while 3 have major defects. What is the probability that a randomly selected unit from the batch has major defects given that it is defective? 0.33 0.25 0.24 0.08

3 A critical system component has a power source consisting of 2 generators only one of which is required to keep the component functional. What is the reliability of this power source if each generator has a reliability of 0.9? 1.0 0.99 0.95 0.9
Answer #1

1.

Correct option: True

2.

Correct option:

0.33

EXPLANATION:

P(Major Defects/Defective) =

3/25 3 9/25- = 0.33 2 = = 0.53 25 9

3.

Correct option:

0.99

EXPLANATION:

Reliability = 1 - (1-0.9)2 = 1 - 0.01 = 0.99